OpenStudy (mhchen):

(MEDAL + FAN) You have 7 males, and 9 females. How many ways can 5 people be selected if at least 2 of them must be male?

OpenStudy (mhchen):

@zepdrix @jigglypuff314 Sorry for tagging, I normally don't, but I don't think many people know how to do this.

jigglypuff314 (jigglypuff314):

I haven't done a question like this before, but I found this similar question that has a solution: http://prntscr.com/cnwwkf

OpenStudy (mhchen):

Very similar, yes.

OpenStudy (mhchen):

From what I got from that example. This is how I attempted the problem: So for the first one, we need 2 male from 7 males, and 3 females from 9 females. etc. \[\left(\begin{matrix}7 \\ 2\end{matrix}\right)\left(\begin{matrix}9 \\ 3\end{matrix}\right) + \left(\begin{matrix}7 \\ 3\end{matrix}\right)\left(\begin{matrix}9 \\ 2\end{matrix}\right)+\left(\begin{matrix}7 \\ 4\end{matrix}\right)\left(\begin{matrix}9 \\ 1\end{matrix}\right)+\left(\begin{matrix}7 \\ 5\end{matrix}\right)\left(\begin{matrix}9 \\ 0\end{matrix}\right)\]

jigglypuff314 (jigglypuff314):

we don't need to consider the 9C0 (last one) but yeah, that looks right ^_^